3. Find the optimal level of production for a monopoly that produces two outputs, x1 and x2, with the following inverse demand functions and cost function: p1 = 88.57 – 0.71 x1 + 0.14 x2 p2 = 77.14 – 0.29 x1 + 0.43 x2 C = 50 – 10 x1 + 20 x2 x1* = ___________________ x2* = ____________________
Answer #1

3 Profit we know that monopolists manimising condition ist MR= MC Now þ= 88.57-0.71 81 +0.1482 Total revenue for product 1 TEquating Mar=mc we get in ... 77,14 -0.29 X1 +0.86 X2 = 2081 029 XL - 14.14X2 =-77.14 0.29x1 + 19.14 X2 = 77,14 -44) Now, sol